Q5) A plane, traveling with a velocity relative to the air of 320 km/h [28° S of W], passes over Winnipeg. The wind
velocity is 72 km/h [S]. Determine the displacement of the plane from Winnipeg 2.0 h later.
1
Expert's answer
2020-10-19T13:21:04-0400
"\bold{V=v_p+v_w}\\=(320\cos{28},320\sin{28})+(0,72)\\=(282.5,222.2)\frac{km}{h}"

"s_x=282.5(2)=565\ km\\s_y=222.2(2)=444.4\ km"

The displacement is


"s=\sqrt{565^2+444.4^2}=720\ km"

"\theta=\arctan{\frac{444.4}{565}}=38\degree\text{ S of W}"